for the pair of functions, find the indicated sum, difference, product, or quotient,
f(x)=x-4, g(x)=-4x^2+16x-4
find (fg)(3)

Answer by jim_thompson5910(35256)

Multiply the two functions (you don't have to algebraically expand everything out) and plug in x = 3 to get...


(f*g)(x) = f(x)*g(x)
(f*g)(x) = ( f(x) )* ( g(x) )
(f*g)(x) = ( x-4 )* ( -4x^2+16x-4 )
(f*g)(3) = ( 3-4 )* ( -4(3)^2+16(3)-4 )
(f*g)(3) = -8